ind Brian's interest rate on a loan if he paid interest of $250 on a principal of $2,000 for a time period of 2 years

Answer by rfer(16322)

I=Prt
r=I/Pt
r=250/(2000*2)
r=250/4000
r=0.0625
r=6.25%